Write down about broadening.
1
Expert's answer
2020-11-12T09:34:56-0500

Solution

When we increase resolution then seperation in energy levels is found which is called broadening.

The relative broadening is written as

"\frac{\Delta E}{E}=\frac{\Delta \nu}{\nu}=\frac{\Delta \lambda }{\lambda}=\frac{\Delta \bar v}{\bar v}"